OpenStudy (anonymous):

plz help .... Find all the eighth roots of (19 + 7 i)

OpenStudy (raden):

eighth roots ?

OpenStudy (anonymous):

means all the roots starting from 0 to 7.... each and every root

terenzreignz (terenzreignz):

I don't that's what it means... I believe you're just tasked to find the eighth roots, and there are many of them... Start by converting it into polar form, it's simple from there on...

OpenStudy (raden):

maybe u want to sum series of (19+7i) from 0 to 7 ??

terenzreignz (terenzreignz):

Well, then it all depends what you mean by 19 + 7i... It sure looks like a complex number to me :D

OpenStudy (anonymous):

terenzreignz u mean (cos @ + i sin@) .....

OpenStudy (anonymous):

yes it is

terenzreignz (terenzreignz):

Yes, I do mean that, but you forgot the "r" it's r cis @

terenzreignz (terenzreignz):

Where cis @ = cos@ + isin@

OpenStudy (anonymous):

like r^(19/7) (cos 19 + i sin 7)

terenzreignz (terenzreignz):

No, that's not how it works, I'm afraid :D For your complex number a + bi can be written in polar form: \[r(\cos \theta + i \sin \theta)\] Where \[r = \sqrt{a^{2}+b^{2}}\] and \[\theta = \tan^{-1}\frac{b}{a}\](not as simple as it seems)

OpenStudy (anonymous):

ok then let me try this one ..... by the way tnx for the help.. :)

terenzreignz (terenzreignz):

Hold, on, before you go, keep this in mind, and think on it deeply... \[\left[ r(\cos \theta + i \sin \theta) \right]^{p}=r^{p}(\cos p\theta + i \sin p\theta)\]

OpenStudy (anonymous):

My θ is 20.22 and r=sqrt410 ..... ?

OpenStudy (anonymous):

correct use radian for your theta=20.22 degrees